OpenStudy (anonymous):

How would i do this? For a certain credit card, the total amount, A ,that a customer owes is given by the following formula. In the formula, P is the starting balance and n is the number of months of debt, assuming interest is compounded monthly. Suppose you begin with a credit card debt of $1000. After one year, you will owe $ _______. A=P(1.015)^n

OpenStudy (anonymous):

dang the credit card companies... they really rip you off eh?

OpenStudy (anonymous):

Given the formula, A=P(1.015)^n, You just need to substitute in the relevant values. P, starting balance, = 1000 n, number of months of debt = one yr = 12 months. A= (1000)(1.015)^12 A= 1195.6 dollars.

OpenStudy (anonymous):

@deszbra